Assume that we are observing a Poisson process in which the average time between events (request for service) is 12 minutes.  Use Excel's EXPONDIST function to compute the following probabilities for the length of time that passes between events:
a.       P(More than 1.5 hours elapse between requests for service)
b.      P(Between 10 and 15 minutes pass between requests for service)
c.       P(Less than 12 minutes elapse)
Answer #1

a)

P(More than 1.5 hours=90 minutes elapse between requests for service) =1-expondist(90,1/12,true)=0.000553

b)

P(10<X<15)=expondist(15,1/12,true)-expondist(10,1/12,true)=0.148093

c)

P( Less than 12 minutes elapse) =expondist(12,1/12,true)=0.632121